# Physics Encoded Blocks in Residual Neural Network Architectures for Digital Twin Models

### v1

@@ Abstract
Physics Informed Machine Learning has emerged as a popular approach in modelling and simulation for digital twins to generate accurate models of processes and behaviours of real-world systems. However, despite their success in generating accurate and reliable models, the existing methods either use simple regularizations in loss functions to offer limited physics integration or are too specific in architectural definitions to be generalized to a wide variety of physical systems. This paper presents a generic approach based on a novel physics-encoded residual neural network architecture to combine data-driven and physics-based analytical models to address these limitations. Our method combines physics blocks as mathematical operators from physics-based models with learning blocks comprising feed-forward layers. Intermediate residual blocks are incorporated for stable gradient flow as they train on physical system observation data. This way, the model learns to comply with the geometric and kinematic aspects of the physical system. Compared to conventional neural network-based methods, our method improves generalizability with substantially low data requirements and model complexity in terms of parameters, especially in scenarios where prior physics knowledge is either elementary or incomplete. We investigate our approach in two application domains. The first is a basic robotic motion model using Euler Lagrangian equations of motion as physics prior. The second application is a complex scenario of a steering model for a self-driving vehicle in a simulation. In both applications, our method outperforms both conventional neural network based approaches as-well as state-of-the-art Physics Informed Machine Learning methods.
@@ 1 Introduction
In this paper, we present two concrete applications to evaluate our method. The first application is based on simulating two-degrees of freedom for a robotic motion as it navigates across specified trajectories. For this, we extend on the work by Lutter et al. where they propose a deep neural network encoding Lagrangian Mechanics (DeLaN) to model the equations of motion for the robotic motion [19]. We formulate their architecture in the context of ’knowledge blocks’ and add residual layers to enhance the learning. For the second application, we test our approach on a more complex real-world problem where we model the behaviour of dynamic steering for autonomous grounded vehicles using a hybrid neural network architecture with the same concept of knowledge blocks. The well-established pure pursuit algorithm provides a physics-derived steering model accounting for dynamics like turn radius. Our architecture includes blocks of fully connected layers to discover intermediate unknown variables in the driving environment. It combines them with differentiable geometrical and kinematic operators from the pure pursuit model. We train this hybrid model observation data from an existing expert driving system inside a simulation environment. The goal is to learn the expert driver’s behaviour or policy using supervised learning. This forms a Behaviour Cloning problem, which involves transforming an expert’s demonstrations into i.i.d state-action pairs and learning to imitate the expert actions given each recorded state using a policy function. In the context of digital twins specifically, this approach is helpful since the underlying goal is to model the exact behaviour of a real-world system using observation data.
This paper presents the following key contributions of our work:
@@
  • It is a significantly more interpretable hybrid neural network comprising learning blocks predicting human-understandable variables and physics blocks predicting the actual target variables.
@@
  • It outperforms the current state-of-the-art Physics Informed Machine Learning model for the Lagrangian-based robotic motion model in terms of accuracy of predicted torque and forces of motion.
  • It outperforms conventional neural networks regarding key driving metrics on unseen road scenarios inside the simulation environment.
  • It uses significantly fewer observation data, with fewer parameters, than conventional neural networks while outperforming key driving metrics inside the simulation environment.
@@ 3 Knowledge Blocks with Residual Learning
4 Experimentation
We validate the effectiveness of our approach in two different experimental setups. The first involves modelling the dynamics of robotic motion using Lagrangian Mechanics (Euler-Lagrange equation) as the Physics Model in our physics-encoded architecture defined earlier. The second setup involves modelling steering dynamics for autonomous vehicle simulation. Here, we investigate the effectiveness of our approach in a more complex digital twin problem where prior physics only partially defines the system’s dynamics.
4.1 Learning Lagrangian Mechanics for Robotic Motion Simulation
We first investigate the effectiveness of our approach, especially the residual architecture’s effect, on a Lagrangian mechanics problem. In this section, we propose to learn the dynamics of robotic motion using the Euler-Lagrange differential equation from Lagrangian mechanics as prior physics knowledge for a learning model. We build on the Deep Lagrangian Networks (DeLaN) by Lutter et al. [19] and extend their architecture in the context of our framework using knowledge blocks. While the authors of DeLaN propose a framework very similar to ours, its applicability is restricted to Lagrangian Mechanics, not addressing the challenges associated with generalizability such as restrictions to gradient flow due to physics-based operators in the neural architecture. To address this, we propose to extend DeLaN architecture with a residual block and present the effectiveness of our framework.
DeLaN incorporates non-conservative forces, generalized coordinates q and generalized velocities $\dot{\textbf{q}}$ , which are the time derivatives of q. Generalized coordinates q refer to coordinates which uniquely identify the configuration of the system being modelled. DeLaN defines the Lagrangian, $L$ as a function of q and $\dot{\textbf{q}}$, as:
$L(\textbf{q},\dot{\textbf{q}})=T(\textbf{q},\dot{\textbf{q}})-V(\textbf{q})$
where $T$ is the kinetic energy and $V$ is the potential energy of the system. For systems undergoing rigid body motion, the kinetic energy can be calculated as the $T=\frac{1}{2}\dot{\textbf{q}}^{\top}\textbf{H}(\textbf{q})\dot{\textbf{q}}$, where $H(\textbf{q})$ is the mass matrix, which is symmetric and positive definite. The Euler-Lagrange equations of motion are given by:
$\frac{d}{dt}\frac{\partial{L}}{\partial{\dot{\textbf{q}}}}-\frac{\partial{L}}{\partial{\textbf{q}}}=\tau$
Hence the Lagrangian equations of motion becomes:
$\textbf{H}(\textbf{q})\ddot{\textbf{q}}+\underbrace{\dot{\textbf{H}}(\textbf{q})\dot{\textbf{q}}-\frac{1}{2}(\frac{\partial}{\partial{\textbf{q}}}(\dot{\textbf{q}}^{\top}\textbf{H}(\textbf{q})\dot{\textbf{q}}))^{\top}}_{\textbf{c}(\textbf{q},\dot{\textbf{q}})}+\textbf{g}(\textbf{q})=\tau$ (1)
[…]
4.1.1 Deep Lagrangian ‘Residual’ Network
We can reformulate the basic architecture behind DeLaN as a combination of knowledge blocks. Since the physics and learning blocks can be conceptually viewed as part of the original DeLaN architecture, we extend the architecture by adding a residual block. Figure 4 illustrates the original architecture from [19] as a combination of learning, residual and physics blocks. The feed-forward layers predicting $\textbf{g}(\textbf{q},\theta)$ and $\textbf{L}(\textbf{q},\theta)$ are illustrated as a learning block that output these intermediate variables into the physics block representing the Euler-Lagrange operators. In the figure, L is divided into off-diagonal ($l_{o}$) and diagonal ($l_{d}$) outputs from two different heads to guarantee positive definiteness. The theoretical understanding behind it can be found in the original paper [19]. The residual block takes as input one of the intermediate layer outputs from the learning block and outputs the residual value $r$. This residual value is added to the motor torque $\tau$ output from the physics block to form the final predicted motor torque value $\hat{\tau}_{target}$.
Figure 4: The computational graph from DeLaN [19] reformulated with learning, residual and physics blocks. The learning block (blue) comprises three separate heads to output g, $l_{o}$ and $l_{d}$. The physics block (orange) represents the mathematical operators from the Euler Lagrange equation. The residual block comprises a set of fully-connected layers, that use an intermediate layer output from the learning block as input. It predicts the residual value $r$ which is added to the predicted torque from the phyics block to output the final torque value $\hat{\tau}_{target}$
The complete computational graph forms our physics-encoded neural network that is trained end-to-end similar to the original DeLaN. The loss function is based on MSE (Mean Squared Error), minimizing the error between the observed ($\tau$ ) and predicted ($\hat{\tau}$) torque values. This is defined as:
$L=\frac{1}{n}\sum_{i=1}^{D}(\tau_{i}-\hat{\tau}_{i})^{2}$
We use the Mini-Batch Gradient Descent algorithm and Adam Optimizer to optimize the loss function, where $n$ is any arbitrary batch size for each sample in the gradient descent algorithm. The gradients for the learnable weights $\theta$ are calculated by the chain of function derivatives that involve the operators from the Euler-Lagrange equation, effectively guiding the learning to follow the physical constraints of the system. The residual blocks facilitate the learning process by providing alternate gradient pathways.
4.2 Steering Dynamics for Autonomous Vehicle Simulation
@@
$\alpha_{axis}$ $[-\pi,+\pi]$ (rad) Angle between the car direction and the direction of the track axis
$D_{center}$ (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) (m) Distance between the car and the track axis. The value is normalized w.r.t to the track width: it is 0 when the car is on the axis, -1 when the car is on the right edge of the track and +1 when it is on the left edge of the car
$z$ [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] (m) Distance of the car mass centre from the surface of the track along the $z$ axis
$V_{x,y,z}$ (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) (km/h) Speed of the car as a vector in $\mathbb{R}^{3}$
$R$ $[0-200]$ (m) Vector of 19 range finder sensors: each sensor returns the distance between the track edge and the car within a range of 200 meters. By default, the sensors sample the space in front of the car every 10 degrees, clockwise from -90 to +90 degrees to the car axis.
$O$ $[0-200]$ (m) Vector of 19 range finder sensors for traffic vehicles: each sensor returns the distance between a detected traffic vehicle and the subject car within a range of 200 meters. By default, the sensors sample the space in front of the car every 10 degrees, clockwise from -90 to +90 degrees to the car axis.
@@
$\delta$ $[-1,+1]$ Steering value: -1 and +1 represent right and left bounds, which corresponds to an angle of 0.366519 rad
$a$ $[0,1]$ Virtual acceleration pedal (0 means no gas, 1 full gas)
$b$ $[0,1]$ Virtual braking pedal (0 means no brake, 1 full brake)
@@ 4.2.1 Physics-based Pure Pursuit Steering Model
Firstly, the algorithm estimates the position of the look-ahead point by projecting it along the desired path based on the vehicle’s current position. Subsequently, the algorithm computes the curvature of the path at the look-ahead point using mathematical techniques such as interpolation or curve fitting. Finally, utilizing the vehicle’s kinematic model, the algorithm calculates the required steering angle to navigate towards the look-ahead point. The look-ahead distance is a control parameter in the context of the algorithm and affects the outcome in two contexts: 1) Regaining a path, i.e., the vehicle is at a “large” distance from the path and must attain the path. 2) Maintaining the path, i.e., the vehicle is on the path and wants to remain on the path. The effects of changing the parameter in the first problem are easy to imagine using the analogy to human driving. The longer look-ahead distances converge to the path more gradually and with less oscillation, as illustrated in Figure 6. It, therefore, depends on the state of the driving environment at any point in time to optimally set the look-ahead distance to manoeuvre the track safely.
@@ 4.2.2 Physics Encoded Residual Feed-Forward Network
The loss function is based on MAE (Mean Absolute Error), minimizing the error between the observed ($\delta$ ) and predicted ($\hat{\delta}$) steering angles. This is defined as:
$L=\frac{1}{n}\sum_{i=1}^{D}|\delta_{i}-\hat{\delta}_{i}|$
We use the Mini-Batch Gradient Descent algorithm and Adam Optimizer to optimize the loss function, where $n$ is any arbitrary batch size for each sample in the gradient descent algorithm. The gradients for the learnable weights in $\phi_{learn}$ are calculated by the chain of function derivatives that involve the geometric and kinematic operators from Equation 2. This allows our feed-forward network to learn the optimal lookahead while being forced to comply with the geometric and kinematic aspects of the steering problem dictated by the pure pursuit algorithm. This is done by constraining the calculation of gradients, which are calculated as an explicit part of the equations based on the pure pursuit algorithm. On the other hand, the residual block provides an alternative pathway for the gradients to flow back to the intermediate layers of the learning block. This allows the loss function to converge and optimise the weights in the residual and, more importantly, the learning block.
@@ 4.2.3 Training on Demonstration Data
Phase II: Train Complete Model
@@
5 Results
In this section, we present experiments and analyses related to the two applications discussed in the last two sections. The first set of experiments related to robotic motion modelling will exhibit the specific improvements that the residual aspect of our approach presents, especially when combining physics and learning blocks in a single neural network architecture. The second set of experiments expands on the broader advantages of this method in achieving higher generalizability with less data and lower model complexity compared to other methods.
5.1 Robotic Motion Simulation
As a base evaluation, we compare our Physics Encoded Residual Neural Network (PERNN) approach to the DeLaN method for modelling Lagrangian dynamics in simulated robotic motion. We use the original experimental setup from [19] for simulated 2-dof robot and compare the effectiveness of our architecture against DeLaN. For experimental consistency, the architecture for the learning and physics blocks are kept constant for both approaches. This presents a fair analysis of the impact of the residual blocks in the optimization process. Since the original work on DeLaN proves the superiority of the base architecture against traditional methods such as feed-forward neural networks, we only compare the results of PERNN with DeLaN to showcase the improvements resulting from our contribution.
5.1.1 Data Generation
Figure 10 illustrates the experimental setup, as presented in [19]. The original work presented three experimental setups: a simulated 2-dof robot drawing the cosine trajectories, a simulated Barrett WAM drawing the 3d cosine 0 trajectory and finally, the physical 7-dof Barrett WAM robot. We limit the scope of experiments to only the simulated 2-dof robot. Here, the generalized coordinates and joint torques are used as input to simulate the character and cosine trajectories using- PyBullet simulation. Table 2 presents the experiment details. 17 unique character trajectories were used as a training set for both models, and the results were calculated and compared over three character trajectories. The training dataset $\mathcal{D}$ consists of state-action pairs as follows:
$\mathcal{D}=\{((\textbf{q}_{1},\dot{\textbf{q}}_{1},\ddot{\textbf{q}}_{1}),\tau_{1}),((\textbf{q}_{2},\dot{\textbf{q}}_{2},\ddot{\textbf{q}}_{2}),\tau_{2}),...((\textbf{q}_{1},\dot{\textbf{q}}_{n},\ddot{\textbf{q}}_{n}),\tau_{n})\}$
Here, q, $\dot{\textbf{q}}$ and $\ddot{\textbf{q}}$ are the generalized coordinates, velocities and acceleration values and $\tau$ is the corresponding motor torque for each instance in time of the simulated robotic joints.
Table 2: Details on the training and test data for modelling 2-dof robotic motion using DeLaN and PERNN methods
Train Samples 2459
Train Characters [’b’, ’o’, ’l’, ’s’, ’u’, ’w’, ’y’, ’a’, ’z’, ’h’, ’m’, ’p’, ’c’, ’d’, ’r’, ’g’, ’n’]
Test Characters [’e’, ’v’, ’q’]
Training Iterations (Epochs) 5000
Figure 10: The experimental setup carried out by [19] (a) The simulated 2-dof robot drawing the character/alphabet trajectories. (b) The simulated Barrett WAM drawing the 3d cosine 0 trajectory. (c) The physical Barrett WAM. This paper presents experiments on (a), a simulated 2-dof robot drawing trajectories of characters/alphabets
5.1.2 Results

### v2

@@ Abstract
Physics Informed Machine Learning has emerged as a popular approach for modeling and simulation in digital twins, enabling the generation of accurate models of processes and behaviors in real-world systems. However, existing methods either rely on simple loss regularizations that offer limited physics integration or employ highly specialized architectures that are difficult to generalize across diverse physical systems. This paper presents a generic approach based on a novel physics-encoded residual neural network (PERNN) architecture that seamlessly combines data-driven and physics-based analytical models to overcome these limitations. Our method integrates differentiable physics blocks—implementing mathematical operators from physics-based models—with feed-forward learning blocks, while intermediate residual blocks ensure stable gradient flow during training. Consequently, the model naturally adheres to the underlying physical principles even when prior physics knowledge is incomplete, thereby improving generalizability with low data requirements and reduced model complexity. We investigate our approach in two application domains. The first is a steering model for autonomous vehicles in a simulation environment, and the second is a digital twin for climate modeling using an ordinary differential equation (ODE)-based model of Net Ecosystem Exchange (NEE) to enable gap-filling in flux tower data. In both cases, our method outperforms conventional neural network approaches as well as state-of-the-art Physics Informed Machine Learning methods.
@@ 1 Introduction
In this paper, we present two concrete applications to evaluate our method. The first application focuses on modeling dynamic steering for autonomous grounded vehicles. In this case, we develop a hybrid neural network architecture that integrates differentiable geometrical and kinematic operators from the well-established pure pursuit algorithm with learning blocks that discover intermediate unknown variables in the driving environment. The model is trained on observation data from an expert driving system in a simulation environment to learn the expert’s behavior or policy using supervised learning. This forms a Behaviour Cloning problem, which involves transforming expert demonstrations into i.i.d. state-action pairs and learning to imitate the expert actions given each recorded state. In the context of digital twins, this approach is particularly useful because it models the exact behavior of a real-world system using observational data.
The second application demonstrates our approach in a real-world digital twin setting for climate modeling. Net Ecosystem Exchange (NEE), a key metric representing the net CO2 exchange between ecosystems and the atmosphere, is typically measured by flux towers that often suffer from data gaps due to instrumentation failures and other issues. To address this challenge, we model NEE using an ordinary differential equation (ODE) derived from an Arrhenius-type formulation of ecosystem respiration. Our Physics Encoded Residual Neural Network (PERNN) for NEE explicitly incorporates the NEE ODE into the network via a dedicated physics block, while learning and residual blocks predict critical intermediate variables such as temperature sensitivity ($E_{0}$), base respiration ($r_{b\_night}$), and the rate of change of ambient air temperature ($\frac{dT}{dt}$). This integration enables accurate gap-filling and forecasting of NEE, leading to improved digital twin fidelity for climate monitoring and forecasting.
This paper presents the following key contributions:
@@
  • It is a significantly more interpretable hybrid neural network comprising learning blocks that predict human-understandable variables and physics blocks that predict the actual target variables.
@@
  • It outperforms conventional neural networks regarding key driving metrics on unseen road scenarios in simulation environments.
  • It uses significantly fewer observation data, with fewer parameters, than conventional neural networks while outperforming key driving metrics.
  • It achieves superior gap-filling performance in digital twin applications for climate modeling by accurately forecasting Net Ecosystem Exchange (NEE) from flux tower data, outperforming state-of-the-art methods such as Random Forest and XGBoost in key error metrics.
  • It provides a robust digital twin framework for environmental monitoring by integrating an ODE-based physics model of ecosystem respiration directly into the neural network architecture.
@@ 3 Knowledge Blocks with Residual Learning
4 Experimental Study
We validate the effectiveness of our approach in two distinct experimental setups. The first study involves modeling steering dynamics for autonomous vehicle simulation. In this scenario, we investigate our method within a simulated digital twin problem where the available prior physics is relatively simple and only partially defines the system’s dynamics. The second study addresses a more complex, real-world digital twin problem for modeling CO2 emissions and gap-filling of Net Ecosystem Exchange (NEE) in flux data. Here, we examine the ability of our approach to integrate an ordinary differential equation that partially characterizes the behavior of emissions with data-driven learning of the remaining dynamics from flux measurements, thereby enabling a forecasting model for gap-filling in emissions data—an important application in the climate digital twin domain
4.1 Case Study 1: Steering Dynamics for Autonomous Vehicle Simulation
@@
$\alpha_{axis}$ $[-\pi,+\pi]$ (rad) Angle between the car direction and the track axis.
$D_{center}$ (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) (m) Normalized distance from the car to the track axis (0 on axis, -1 at the right edge, +1 at the left edge).
$z$ (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) (m) Vertical distance of the car’s mass centre from the track surface.
$V_{x,y,z}$ (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) (km/h) Speed vector in $\mathbb{R}^{3}$.
$R$ $[0,200]$ (m) 19 range finder sensors: each returns the distance from the car to the track edge, sampling every 10° from -90° to +90° relative to the car’s axis.
$O$ $[0,200]$ (m) 19 range finder sensors for traffic vehicles with the same angular sampling.
@@
$\delta$ $[-1,+1]$ Steering value: -1 and +1 represent the right and left bounds, corresponding to 0.366519 rad.
$a$ $[0,1]$ Virtual acceleration pedal (0 = no gas, 1 = full gas).
$b$ $[0,1]$ Virtual braking pedal (0 = no brake, 1 = full brake).
@@ 4.1.1 Physics-based Pure Pursuit Steering Model
Firstly, the algorithm estimates the position of the look-ahead point by projecting it along the desired path based on the vehicle’s current position. Subsequently, the algorithm computes the curvature of the path at the look-ahead point using mathematical techniques such as interpolation or curve fitting. Finally, utilizing the vehicle’s kinematic model, the algorithm calculates the required steering angle to navigate towards the look-ahead point. The look-ahead distance is a control parameter in the context of the algorithm and affects the outcome in two contexts: 1) Regaining a path, i.e., the vehicle is at a ”large” distance from the path and must attain the path. 2) Maintaining the path, i.e., the vehicle is on the path and wants to remain on the path. The effects of changing the parameter in the first problem are easy to imagine using the analogy to human driving. The longer look-ahead distances converge to the path more gradually and with less oscillation, as illustrated in Figure 5. It, therefore, depends on the state of the driving environment at any point in time to optimally set the look-ahead distance to manoeuvre the track safely.
@@ 4.1.2 Physics Encoded Residual Feed-Forward Network
The loss function is based on Mean Squared Error (MSE), minimizing the error between the observed ($\delta$ ) and predicted ($\hat{\delta}$) steering angles. This is defined as:
$L=\frac{1}{n}\sum_{i=1}^{n}(\delta_{i}-\hat{\delta}_{i})^{2}$
We use the Mini-Batch Gradient Descent algorithm and Adam Optimizer to optimize the loss function, where $n$ is any arbitrary batch size for each sample in the gradient descent algorithm. The gradients for the learnable weights in $\phi_{learn}$ are calculated by the chain of function derivatives that involve the geometric and kinematic operators from Equation 1. This allows our feed-forward network to learn the optimal lookahead while being forced to comply with the geometric and kinematic aspects of the steering problem dictated by the pure pursuit algorithm. This is done by constraining the calculation of gradients, which are calculated as an explicit part of the equations based on the pure pursuit algorithm. On the other hand, the residual block provides an alternative pathway for the gradients to flow back to the intermediate layers of the learning block. This allows the loss function to converge and optimize the weights in the residual and, more importantly, the learning block.
@@ 4.1.3 Training on Demonstration Data
Phase II: Integrated Training
@@
4.2 Case Study 2: Climate Digital Twins for CO2 Emissions and Gap-Filling in Net Ecosystem Exchange Modelling
The application of digital twin technology to climate systems has emerged as a promising avenue for enhanced monitoring and forecasting of atmospheric CO2 dynamics. Digital twins, which provide a virtual representation of real-world processes, are particularly valuable in climate science where continuous, high-quality observational data are critical yet often incomplete. Flux towers continuously monitor key atmospheric scalars such as Net Ecosystem Exchange (NEE), Latent Heat, and Sensible Heat [32] using the eddy covariance (EC) method. However, these measurements frequently exhibit gaps due to instrumentation failures, power shortages, or environmental interferences [33, 34].
In the context of a digital twin for climate applications, robust gap-filling methods are essential not only for data completion but also for accurately capturing the underlying physical processes governing NEE. Traditional gap-filling techniques—such as Non-Linear Regression, Look-Up Tables, and Marginal Distribution Sampling—have shown reasonable performance for short-duration gaps, yet they struggle with longer intervals of missing data [34]. More recent approaches, such as the Random Forest Robust (RFR) method, have enhanced performance over longer gaps by improving the statistical robustness of the estimates, although challenges remain for nighttime measurements [32].
Integrating the underlying physics of NEE into the gap-filling process can address these challenges. As demonstrated in previous work, NEE dynamics can be modelled as a Stochastic Differential Equation (SDE) that decomposes the signal into a deterministic drift component—governed by temperature-dependent ecosystem respiration following an Arrhenius-type law [35, 36]—and a stochastic diffusion term representing measurement noise, assumed to be Gaussian [37]. Embedding this SDE within a digital twin framework allows the inherent physics of ecosystem CO2 exchange to directly inform the data reconstruction process, leading to more reliable and interpretable predictions.
Our proposed Physics Encoded Neural Network (PERNN) framework naturally extends to this climate modelling use case. In this setting, the physics blocks encapsulate the SDE-based model of NEE, capturing both the deterministic Arrhenius-type response of ecosystem respiration and the inherent stochastic fluctuations. Complementary learning blocks infer latent variables and compensate for unobserved dynamics during periods of data absence. Moreover, residual blocks not only aid gradient flow during training but also facilitate the integration of these physics-informed estimates with empirical corrections, ensuring that the digital twin maintains fidelity to both the established physical laws and the observational data.
This hybrid architecture enhances gap-filling performance while simultaneously improving the interpretability of NEE trends, making it a robust tool for climate digital twins. By directly incorporating physical models into the neural network, our approach bridges the gap between purely data-driven methods and traditional physics-based modeling, thereby providing a scalable and generalizable framework for environmental monitoring and forecasting [38, 39].
4.2.1 Flux Tower Data
The observational data employed in this study were collected from a flux tower located in East Anglia, UK [40] from 2012 to 2019. The tower is equipped with an eddy covariance (EC) system designed to measure key atmospheric scalars including Net Ecosystem Exchange (NEE), sensible heat flux density (H), momentum flux (Tau), relative humidity (RH), vapor pressure deficit (VPD), global radiation ($\text{R}_{\text{g}}$), friction velocity (Ustar), soil temperatures (Tsoil1 and Tsoil2), and air temperature at 2 m ($\text{T}_{\text{air}}$). The dataset spans eight years, from 2012 through the end of 2019, with measurements recorded at 30-minute intervals.
These high-frequency measurements provide a rich temporal resolution of the ecosystem’s CO2 exchange dynamics. However, the dataset also presents common challenges such as missing values caused by power shortages, sensor malfunctions, or adverse weather conditions. These data gaps can range from short intervals to periods extending over several months, emphasizing the necessity for robust gap-filling methods that can reliably reconstruct missing measurements while preserving the underlying physical relationships.
A detailed description of the measured variables is provided in Table 4.2.1. In this study, particular attention is given to the NEE variable, which is central to understanding the carbon dynamics of the ecosystem. Given that latent heat (L) is measured by the same instrument as NEE, any gaps in NEE often correspond to gaps in latent heat measurements; hence, latent heat is excluded from the analysis to maintain consistency. Additionally, time-based attributes such as season, hour, day of week, month, and day of year are incorporated to account for diurnal and seasonal variations.
Table 2: List of variables from the flux data [40]
Variable Units Description